A small boat has a mass of 400 kg and a total volume of .8m^3. It floats on salt water with a density of 1030kg/m^3.

A) what is the mass of the water displaced by the boat?

B) what is the volume of the water displaced by the boat?

C) A 35kg child gets into the boat. What is the volume of the water displaced now?
Answer #1

part a )

Mdis = mg = 400 * 9.8 = 3920 kg

part b )

Mdis = rho * Vdis

Vdis = 3920/1030 = 3.8 m^3

part c )

Mdis = 4263 kg

Vdis = 4263/1030 = 4.1 m^3